Circular No. 6190

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 O1

S. Nakano, Sumoto, Japan, reports the following additional precise positions:

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 24.58491 18^{h}43^{m}07.^{s}15 −32^{o}08′29′′.1  Kushida
24.69115 18 43 02.59 −32 08 19.8  Kobayashi
24.69782 18 43 02.19 −32 08 18.9  _{”}
25.64873 18 42 21.92 −32 06 45.4  Sugie
25.65515 18 42 21.60 −32 06 44.9  _{”}
25.65862 18 42 21.50 −32 06 44.3  _{”}
25.66279 18 42 21.33 −32 06 44.5  _{”}
26.50591 18 41 45.61 −32 05 20.4  Kojima
26.50833 18 41 45.72 −32 05 20.0  Kushida
26.51528 18 41 45.41 −32 05 19.0  _{”}
26.51852 18 41 45.33 −32 05 20.0  Kojima
26.52222 18 41 45.09 −32 05 18.8  Kushida
26.55608 18 41 43.67 −32 05 16.6 10.4 Nakamura
26.56042 18 41 43.48 −32 05 16.1  _{”}
26.56186 18 41 43.41 −32 05 14.8  Sugie
26.56458 18 41 43.30 −32 05 15.7  Nakamura
26.57101 18 41 43.04 −32 05 14.2  Sugie
26.57486 18 41 42.97 −32 05 11.7  Abe
26.60867 18 41 41.34 −32 05 07.7  _{”}

Y. Kushida (Yatsugatake South Base Observatory). 0.25-m f/6.3 Schmidt-

Cassegrain reflector + CCD. J. Kobayashi (Kumamoto Civil Astronomical Observatory). 0.41-m f/5.0

reflector + CCD. A. Sugie (Dynic Astronomical Observatory). 0.60-m f/5.0 reflector + CCD. T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6.0 reflector + CCD. A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien + CCD. H. Abe (Yatsuka). 0.26-m f/4.8 reflector + CCD.

Total visual magnitude estimate by R. Keen, Mt. Thorodin, CO (0.32m reflector): July 25.24 UT, 10.4 (coma diameter 3′).

1995 July 26 Daniel W. E. Green